Kinkaid Co. is incorporated at the beginning of this year and engages in a number of transactions. The following journal entries impacted its stockholders’ equity during its first year of operations. General Journal Debit Credit a. Cash 290,000 Common Stock, $25 Par Value 250,000 Paid-In Capital in Excess of Par Value, Common Stock 40,000 b. Organization Expenses 200,000 Common Stock, $25 Par Value 128,000 Paid-In Capital in Excess of Par Value, Common Stock 72,000 c. Cash 45,000 Accounts Receivable 17,000 Building 82,200 Notes Payable 59,600 Common Stock, $25 Par Value 54,600 Paid-In Capital in Excess of Par Value, Common Stock 30,000 d. Cash 131,000 Common Stock, $25 Par Value 75,000 Paid-In Capital in Excess of Par Value, Common Stock 56,000 Required: 2. How many shares of common stock are outstanding at year-end? 3. What is the amount of minimum legal capital (based on par value) at year-end? 4. What is the total paid-in capital at year-end? 5. What is the book value per share of the common stock at year-end if total paid-in capital plus retained earnings equals $787,000?
Answer #1

2. Shares O/s at the end of year = a + b + c + d

Shares O/s at the end of year = 250000 / 25 + 128000/25 + 54600 / 25 + 75000 / 25

Shares O/s at the end of year = 10000 + 5120 + 2184 + 3000

Shares O/s at the end of year = 20304 Shares

3. Minimum Legal Capital = Shares O/s * Par Value

Minimum Legal Capital = 20304 Shares * $25

Minimum Legal Capital = $507600

4. Total Paid in capital = $507600 + 40000 + 72000 + 30000 + 56000

Total Paid in capital = $705600

5. Book Value per Share = $787000 / 20304

Book Value per Share = $38.76